While traveling to his grandmother's for Easter, George fell asleep halfway through the journey. When he awoke, he still had to travel half the distance that he had traveled while sleeping. For what part of the entire journey had he been asleep?

Thanks For you Time!
Ian

Answer by checkley71(8403)

1/2+1/x+1/2x=1 or (x+2+1)/2x=1 or 2x=x+2+1 or x=3 thus 1/2+1/3+1/2*3=1 or
he had been sleeping 1/3 of the trip.
1/2+1/3+1/6=1 or 3/6+2/6+1/6=1 or 6/6=1 or 1=1